OpenStudy (anonymous):

I have been having trouble understanding literal equations so I am using an example from my algebra homework...... 3+6x=11-4y

OpenStudy (anonymous):

Are you trying to solve for y?

OpenStudy (anonymous):

yes it says to solve Y as a function of X

OpenStudy (anonymous):

Ok. So to "solve for" a variable means you get it by itself. So first we should get rid of that 11. Remember you have to do the same thing to both sides of the equation. So we subtract 11 from both sides. \[3 + 6x - 11 = 11 - 11 - 4y\]\[6x - 8 = -4y\]

OpenStudy (anonymous):

ok I got the same answer this is always where I get stuck

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

oh nevermind I got it but ty